OpenStudy (anonymous):

Two people, Lesley and Jaime, are racing each other. Assume that both their accelerations are constant, Lesley covers the last 1/6 of the race in 3 seconds, and Jaime covers the last 1/3 of the race in 6 seconds. Who wins, and by how much?

OpenStudy (ash2326):

If we could find their acceleration, then we can find the time it takes for them to finish the race. @vancouver2012 how would relate distance with acceleration and time?

OpenStudy (anonymous):

derivatrive

OpenStudy (ash2326):

\[S=ut+\frac 12 at^2\] Do you know this equation?

OpenStudy (anonymous):

in physics yea

OpenStudy (ash2326):

I'm thinking, how we'd use this.

OpenStudy (anonymous):

t=sqrt(2*d/a)

OpenStudy (ash2326):

Initial velocity=0 for the starting of the race not for the scenario we are given

OpenStudy (anonymous):

the answer key says use t=sqrt(2*d/a)

OpenStudy (anonymous):

its given as a hint

OpenStudy (ash2326):

Lesley covers the last 1/6 of the race in 3 seconds let the race length be x so \[\frac {x}{6}=u1 \times 3+\frac 12 a (3)^2\] u1= initial velocity of Lesley for the last 1/6 of the race Similarly for Jaime, we write this equation \[\frac {x}{3}=u2 \times 6+\frac 12 a (6)^2\] also \[v^2-u^2=2as\] for lesley \[(u1)^2=2a \times \frac56 x\] for Jaime \[(u2)^2=2\times \frac 2 3 x\] do you get this ?

OpenStudy (anonymous):

not this v 2 −u 2 =2as

OpenStudy (anonymous):

also this question is listed in derivative section

OpenStudy (anonymous):

Using this formula, we find that the time (in seconds) it takes Lesley to run the race is and that the time it takes Jaime is

OpenStudy (anonymous):

that is what the hint says after providing the formula that I provided

OpenStudy (ash2326):

for Leslie v=u1 u=0